Circular No. 5656

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

SUPERNOVA 1992bj IN ANONYMOUS GALAXY C. Pollas, Observatoire de la Côte d’Azur, reports his discovery, on

films taken Oct. 17.90 and 18.93 UT, of a supernova (mag ∼ 19.5-20) located ∼ 0′′.1 west and 4′′.9 north of the nucleus of a galaxy of mag ∼ 18 (α = 23^{h}19^{m}21^{s}.8, δ = +24^{o}53′02′′, equinox 1950.0), near SAO 91205. The supernova’s position has end figures 22^{s}.42, 06′′.4. Nothing is visible at this location on a Technical Pan plate from 1990 or on the prints of the Palomar Sky Survey.

A. V. Filippenko, T. Matheson, and L. C. Ho, University of California at Berkeley, report that CCD spectra (range 320-1000 nm) obtained on Nov. 18 UT with the Shane 3-m reflector at Lick Observatory reveal that this object is a supernova of type II. Prominent Hα having a P-Cyg profile is present. The redshift of the parent galaxy is roughly 0.056, measured from the narrow Hα emission of the superposed H II region. Additional photometry and spectroscopy would be useful in order to determine the distance to the supernova with the expanding photosphere method.

HR 1469 A. Bradley, Hubble Space Telescope (HST) Operations, Goddard Space

Flight Center; and D. Hoffleit and I. Platais, Department of Astronomy, Yale University, report: “During the routine operation of the HST Fixed Head Star Trackers on Nov. 4.642 UT, a considerable fade in brightness, to mag 6.9 for at least 3 hr, was noticed of the star HR 1469 = 49 Eri = NSV 1671 (normally at V = 5.31). Cousins (1963, MNASSA 22, 12) reported on possible variability of 0.08 mag, but the large decrease in brightness observed by the FHST has not been recorded before, to our knowledge. To investigate in more detail the nature of the light variability in HR 1469, we would encourage professional and amateur astronomers to monitor for the possible light change as well as to check the photographic plates for other minima in the past.”

NOVA IN THE LARGE MAGELLANIC CLOUD 1992 Visual magnitude estimates: Nov. 13.460 UT, 10.2 (P. Camilleri, Co-

bram, Victoria, Australia); 13.854, 10.7 (J. Hers, Sedgefield, South Africa); 14.078, 10.65 (W. Liller, Viña del Mar, Chile; CCD); 14.597, 10.7 (Camilleri); 14.826, 10.9 (Hers); 15.204, 10.87 (Liller; CCD); 15.451, 11.0 (Camilleri); 16.144, 11.33 (Liller; CCD); 16.475, 11.5 (Camilleri).

1992 November 18 Daniel W. E. Green